OpenStudy (anonymous):

Joey is selling tickets for a high school play. Student tickets cost $4 and general admission tickets cost $6. He sells 525 tickets total. He collects $2,876.00. S=number of student tickets. G=number of general admission tickets. Write a system of equations that can be used to find the number of students tickets, S and the number of general admission tickets, G that Joey sold. Solve the system!

OpenStudy (anonymous):

Can anyone help me with this?

OpenStudy (anonymous):

s + g = 525 4s + 6g = 2876 Multiply the first equation by -4 :)

OpenStudy (anonymous):

-4s -4g = 2100 {-4s - 4g} = 2100 + 4s + 6g = 2876 = 0s + 2g = 776 2g = 776 g = 388 So it's 388 general admission tickets. I'm sure you can do the math to find the student tickets :)

OpenStudy (anonymous):

why is it not -2100??

OpenStudy (anonymous):

It is! Sorry, my keyboard is stuck!

OpenStudy (anonymous):

thought so! no biggie!

OpenStudy (anonymous):

im confused on this line {-4s - 4g} = 2100 + 4s + 6g = 2876 = 0s + 2g = 776

OpenStudy (anonymous):

NVM :)

OpenStudy (anonymous):

{-4s - 4g} = -2100 + 4s + 6g = 2876 You eliminate 4s, subtract 4g from 6g, and subtract 2100 from 2876. (: